# Co-Design and Evaluation of a CPU-Free MPI GPU Communication Abstraction and Implementation

### v1

@@ II-C Slingshot 11/libfabric GPU Communication Features
libfabric defines two features to enable construction of triggered communication: (1) triggering counters that count network device events and can be configured to increment either upon the occurrence of specific network events (e.g., the local completion of a local operation such as a send, atomic operation, a remote write from a remote network device) or by explicit manipulation by the host, and (2) deferred work queue entries (DWQs) that specify libfabric communication operations222libfabric deferred work queue entries currently support only point-to-point communication operations, not hardware collective operations. whose execution is delayed until an associated triggering counter reaches a specified integer value.
@@ III-B MPI_Queue: Ordering Progress with External Execution
To represent an ordered set of communication tasks to execute on a GPU, we define an MPI_Queue object. Proposed APIs related to general management of the Queue object itself can be found in in Table I. These APIs are similar to other creation and freeing functions found within the MPI standard. The remainder of this section explains both how to use the proposed Queue objects.
@@ VI-A2 GPU Ping-Pong Microbenchmark
After completing warm-up iterations, we ran each ping-pong trial for a given buffer size for a number of round-trip iterations chosen to make the total measurement time between between 1 and 30 seconds; this was 100,000 iterations for buffer sizes less than 4MB, 10,000 iterations for buffer sizes between 4MB and 64MB, and 1,000 iterations for buffer sizes greater than 64MB. We ran five ping-pong trials for all power-of-two buffer sizes between one byte and one gigabyte, and computed and plotted the average and 95% confidence interval of the per-trial average ping-pong latency and bandwidth.
@@ Acknowledgements
This research used resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725.

### v2

# Patrick G. Bridges    Derek Schafer    Jack Lange James B. White III    Anthony Skjellum    Evan Suggs Thomas Hines    Purushotham Bangalore    Matthew G. F. Dosanjh    Whit Schonbein

@@ II-C Slingshot 11/libfabric GPU Communication Features
libfabric defines two features to enable construction of triggered communication: (1) triggering counters that count network device events and can be configured to increment either upon the occurrence of specific network events (e.g., the local completion of a local operation such as a send, atomic operation, or a remote write from a remote network device) or by explicit manipulation by the host, and (2) deferred work queue entries (DWQs) that specify libfabric communication operations222libfabric deferred work queue entries currently support only point-to-point communication operations, not hardware collective operations. whose execution is delayed until an associated triggering counter reaches a specified integer value.
@@ III-B MPI_Queue: Ordering Progress with External Execution
To represent an ordered set of communication tasks to execute on a GPU, we define an MPI_Queue object. Proposed APIs related to general management of the Queue object itself can be found in in Table I. These APIs are similar to other creation and freeing functions found within the MPI standard. The remainder of this section explains how to use the proposed Queue objects.
@@ VI-A2 GPU Ping-Pong Microbenchmark
After completing warm-up iterations, we ran each ping-pong trial for a given buffer size for a number of round-trip iterations chosen to make the total measurement time between 1 and 30 seconds; this was 100,000 iterations for buffer sizes less than 4MB, 10,000 iterations for buffer sizes between 4MB and 64MB, and 1,000 iterations for buffer sizes greater than 64MB. We ran five ping-pong trials for all power-of-two buffer sizes between one byte and one gigabyte, and computed and plotted the average and 95% confidence interval of the per-trial average ping-pong latency and bandwidth.
@@ Acknowledgements
This research used resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Advanced Scientific Computing Research programs in the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725.